# Ground states of strongly-correlated materials on quantum computers using ab initio downfolding

### v1

@@ Abstract
The accurate first-principles description of strongly-correlated materials is an important and challenging problem in condensed matter physics. Ab initio downfolding has emerged as a way of deriving accurate many-body Hamiltonians including strong correlations, representing a subspace of interest of a material, using density functional theory calculations as a starting point. However, the solution of these material-specific models can scale exponentially on classical computers, constituting a challenge. Here we propose that utilizing quantum computers for obtaining the properties of downfolded Hamiltonians yields an accurate description of the ground state properties of strongly-correlated systems, while circumventing the exponential scaling problem. We demonstrate this for diverse strongly-correlated materials by combining ab initio downfolding and variational quantum eigensolvers, correctly predicting the antiferromagnetic state of one-dimensional cuprate $\text{Ca}_{2}\text{CuO}_{3}$, the excitonic ground state of monolayer $\text{WTe}_{2}$, and the charge-ordered state of correlated metal $\text{SrVO}_{3}$. By utilizing a classical tensor network implementation of variational quantum eigensolvers we are able to simulate large models with up to $54$ qubits and encompassing up to four bands in the correlated subspace, which is indicative of the complexity that our framework can address.
@@ I Introduction
A promising computational method for describing strongly-correlated systems from first principles is ab initio downfolding [11, 12, 13, 14], which utilizes a DFT starting point in order to derive a material-specific many-body Hamiltonian within a strongly-correlated active space of interest, which can be subsequently solved using a range of techniques, including exact diagonalization [15, 16], quantum Monte Carlo [17], diagrammatic approximations [18], and beyond. The many-body Hamiltonian resulting from downfolding and describing the physics of the strongly-correlated region is typically of the extended Hubbard form [19]:
@@
where $\sigma,\rho$ are spin indices, $\mathbf{R},\mathbf{R}^{\prime}$ denote lattice vectors, and $i,j$ run over the electronic bands of the system. Downfolding methods have been successfully applied to the description of charge-ordered systems and charge density waves [20, 21], high-temperature superconductors [13, 22, 23, 24, 25], and beyond. However, the unfavorable scaling of classical methods for obtaining eigenstates of the Hamiltonian of Eq. (I) has thus far prevented downfolding methods from being used for large systems, which can result in finite-size effects [26, 27], and limits the number of bands which may realistically be included in these models.
@@
Quantum computers are a promising technology for the simulation of many-body quantum systems [28, 29], with a theoretical linear scaling of the computational cost with the number of orbitals, in an idealized scenario, given the 1:1 correspondence between qubits and spin-orbitals in models such as that of Eq. (I). This constitutes the simulation of strongly-correlated systems and Hubbard-like models an ideal problem for solution on quantum hardware [30, 31], and the performance of popular variational quantum algorithms in terms of producing an accurate ground state has recently been the subject of a detailed benchmark [32]. The potential of emerging fault-tolerant quantum computers to obtain properties of extended Hubbard models as that of Eq. (I) has been discussed in Ref. [33], where the necessary resources to access experimentally-relevant quantities have been estimated.
@@
Here we demonstrate that current and near-term quantum computers can produce quantitatively and qualitatively accurate results on the strongly-correlated ground states of diverse materials. We do so by combining ab initio downfolding, in order to prepare material-specific Hamiltonians of the form of Eq. (I), with classical simulations of variational quantum eigensolver (VQE) optimizations to obtain the ground state of these Hamiltonians, utilizing a recently proposed tensor network implementation [34]. We show that our VQE energy for the ground state is quantitatively accurate when compared to results obtained within density matrix renormalization group (DMRG), and that we obtain the correct behavior for the ground state wavefunction and correlation functions across different scenarios, and specifically the antiferromagnetic behavior of the quasi-1D cuprate $\text{Ca}_{2}\text{CuO}_{3}$ [35], the excitonic insulating ground state of the two-dimensional material $\text{WTe}_{2}$ [36, 7], and the charge-ordered state of the correlated metal $\text{SrVO}_{3}$ [37, 38]. These results, combined with the demonstrations that quantum simulation of downfolded Hamiltonians can yield accurate excited state properties in strongly correlated molecular systems [39, 40, 41], demonstrate the strong potential of current and emerging quantum computers for simulating the properties of strongly-correlated materials when combined with ab initio downfolding.
@@ II.1 Ab initio downfolding
The aim of ab initio downfolding approaches is to generate a many-body Hamiltonian representing an active space of interest for a given material, typically within the low-energy region in the vicinity of the Fermi level [11, 12, 13, 19], most often starting from a first-principles description such as DFT. An exchange term may be included in downfolded Hamiltonians [19], however we will ignore it here given its small magnitude for our studied systems, consistent with previous works [20]. The Hamiltonian of Eq. (I) obtained through ab initio downfolding contains hopping and Coulomb terms between all neighbors on a lattice, and indeed some of the longer-range terms, particularly next-nearest neighbor coupling, can be important in certain cases [42, 43]. Here we include nearest-neighbor terms only for the hopping and Coulomb terms, as these dominate over longer-range terms in the Wannier representation, due to the exponential decay of Wannier functions [44].
In order to describe periodic solids as a lattice model, we work within the basis of maximally-localized Wannier functions [45], which we will denote as $\phi$. Here the Wannier functions are centered at the different lattice sites $\mathbf{R}$ appearing in Eq. (I), with $i,j$ indexing the $i^{\text{th}}$ and $j^{\text{th}}$ Wannier function respectively. Starting from a Kohn-Sham DFT calculation performed within Quantum Espresso [46], the Wannier representation of the studied systems within the active space of interest is obtained using the wannier90 code [47]. The Wannierization procedure yields the hopping terms appearing in Eq. (I) as
@@
where $H_{KS}$ the Kohn-Sham Hamiltonian and $V$ the volume of the unit cell. For $\mathbf{R}=\mathbf{R}^{\prime}$ this term represents the chemical potential of site $\mathbf{R}$. Moreover, following the wannierization using wannier90, we utilize the wan2respack [48] and RESPACK [19] software packages in order to obtain the two-body Coulomb integrals appearing in Eq. (I) as
@@ II.2 Tensor network VQE simulation
Our aim is to obtain the ground state of the Hamiltonian of Eq. (I), for studied materials with an active space represented by different sets of $U_{i\mathbf{R}j\mathbf{R}^{\prime}}$ and $t_{i\mathbf{R}j\mathbf{R}^{\prime}}$ matrices. Before presenting our approach for computing the ground state, it is worth discussing some aspects of the complexity of this problem. For a Hubbard model on an $N_{x}\times N_{y}$ square lattice, with $N_{b}$ electronic bands included in the active space of the studied system, $n_{q}=2N_{x}N_{y}N_{b}$ is the number of qubits needed to simulate its properties, where the factor $2$ accounts for spin. Denoting $N=N_{x}N_{y}$, the number of terms appearing in the Hamiltonian of Eq. (I) is classified as follows, considering only nearest-neighbor contributions:
  • Chemical potential $t_{i\mathbf{R}i\mathbf{R}}$: For the on-site energies, there are $2\times N_{b}\times N$ terms.
  • Total number of terms:
    $N_{b}\times\bigg{[}(N_{b}+2)[N_{x}\times(N_{y}-1)+$
    $N_{y}\times(N_{x}-1)]+3N_{b}N\bigg{]}$
We do not include inter-band hopping between different sites to ensure a manageable computational cost for the systems studied here. Boundary effects have been implicitly considered by restricting hopping and off-site interactions to existing nearest-neighbor pairs within the lattice.
@@
Unless otherwise explicitly stated, we use the maximum bond dimension $\chi_{\rm max}=2^{n_{q}/2}$, where $n_{q}$ the number of qubits. Since this bond dimension is sufficient to exactly represent an arbitrary wavefunction on $n_{q}$ qubits, the MPS representation yields the energy expectation value with respect to the exact PQC. For all tensor operations in this work we have used the ITensor software package [53]. We provide details on our optimization procedure in Section S1 of the Supplementary Information (SI). In preparing our PQC, we utilize two ansätze in this work, a number-preserving (NP) ansatz [31], which was designed specifically within the context of solving single-band Hubbard models, and a more generic excitation-preserving (EP) [54] ansatz, which allows us to straightforwardly prepare our circuits in the case of multi-band Hubbard models. The performance of both ansätze for producing accurate ground states has been discussed in Ref. [32]. We apply the ansätze to the solution of the non-interacting ($U=0$) Hubbard representation of our systems, prior to which we apply $R_{z}(\theta)$ gates to each qubit, which improves the optimization procedure.
@@
It is worth emphasizing that the quantum simulation of materials using representations as the Hamiltonian of Eq. (I) has been the subject of detailed benchmarks, and found to likely be feasible with near-term resources, without necessitating fully fault-tolerant quantum computers [55]. Nevertheless, as we aim to build increasing complexity into the reduced Hamiltonians obtained through downfolding or related methods, and approach the full many-body limit, it might become necessary to utilize fault-tolerant architectures [33]. However, our work here is mostly agnostic to this.
@@
With regards to the fault tolerant era, we perform a rough estimate of the resources needed for the ansatz considered here, and for this purpose it is interesting to estimate the number of $T$-gates which will be necessary in order to implement our circuits in a fault-tolerant fashion, where the one-qubit and two-qubit operators are approximated with a required accuracy $\epsilon$. As described previously, we apply an $R_{z}$ gate to each qubit prior to the application of the unitary operators associated with the variational ansatz. The approximation of a one-qubit gate has been found to be possible with $1.15\log_{2}(1/\epsilon)$ $T$-gates [56]. From this we have $n_{q}\times 1.15\log_{2}(1/\epsilon)$ $T$-gates associated with this step. Additionally, both the NP and EP ansätze involve the application of two-qubit gates with two parameters each. For a total of $n_{\text{params}}$ parameters in the variational ansatz we have $n_{\text{params}}/2$ two-qubit gates, each of which requires $\mathcal{O}[16\log(1/\epsilon)+32]$ $T$-gates to represent [57]. Therefore, we have a total of $n_{q}\times 1.15\log_{2}(1/\epsilon)n+\frac{n_{\text{params}}}{2}\times\mathcal{O}[16\log(1/\epsilon)+32]$ $T$-gates required for the fault-tolerant simulation of our systems with accuracy $\epsilon$.
@@ III Results
$\text{Ca}_{2}\text{CuO}_{3}$ $10\times 1$ $1$ $-1.030$ $-1.019$ $99.8\%$
$\text{WTe}_{2}$ $2\times 2$ $4$ $55.103$ $55.225$ $90.5\%$
$\text{SrVO}_{3}$ $3\times 3$ $3$ $-116.192$ $-115.995$ $76.8\%$
Table 1: Studied systems, simulated lattice sizes, number of bands in the subspace, DMRG energies, and best values for the VQE energies and fidelities with respect to the DMRG solution. As outlined in the main text, we solve the Hubbard model representation of $\text{Ca}_{2}\text{CuO}_{3}$ at half-filling, of $\text{WTe}_{2}$ at full filling of the valence bands, and of $\text{SrVO}_{3}$ at half-filling of the lowest band.
Fig. 1 visualizes the structures of the systems we study here. These are chosen as demonstrate the capability of our approach to correctly predict the ground state properties of diverse strongly-correlated materials; $\text{Ca}_{2}\text{CuO}_{3}$ is a quasi-1D cuprate known to display antiferromagnetic behavior along the Cu-O chains [35]; $\text{WTe}_{2}$ has been proposed to host an excitonic ground state, i.e., one where correlated electron-hole pairs form spontaneously [36, 7]; and $\text{SrVO}_{3}$ is a correlated metal that exhibits substantial charge ordering [37, 38]. Table 1 summarizes the number of bands included in the active space of each system, the lattice size on which the downfolded Hamiltonian is solved, the DMRG energy, and the best value for the energy and fidelity obtained from the ten independent VQE optimizations we perform. For $\text{Ca}_{2}\text{CuO}_{3}$ we only perform a single optimization, as the simpler energy landscape of this system makes this sufficient for finding a low-energy, high-fidelity solution. More details on the optimization for each system are given below. In Section S1 of the SI we also give the VQE energy and fidelity values for all optimizations performed for $\text{WTe}_{2}$ and $\text{SrVO}_{3}$. Moreover, we summarize details of all DFT and cRPA calculations employed for these systems in SI Section S2. We now discuss the properties of the ground state wavefunctions of the materials of Fig. 1 as obtained within our combined ab initio downfolding/VQE approach.

### v3

# Compressing Hamiltonians with ab initio downfolding for simulating strongly-correlated materials on quantum computers

@@ Abstract
The accurate first-principles description of strongly-correlated materials is an important and challenging problem in condensed matter physics. Ab initio downfolding has emerged as a way of deriving compressed many-body Hamiltonians that maintain the essential physics of strongly-correlated materials. The solution of these material-specific models is still exponentially difficult to generate on classical computers, but quantum algorithms allow for a significant speed-up in obtaining the ground states of these compressed Hamiltonians. Here we demonstrate that utilizing quantum algorithms for obtaining the properties of downfolded Hamiltonians can indeed yield high-fidelity solutions. By combining ab initio downfolding and variational quantum eigensolvers, we correctly predict the antiferromagnetic state of one-dimensional cuprate $\text{Ca}_{2}\text{CuO}_{3}$, the excitonic ground state of monolayer $\text{WTe}_{2}$, and the charge-ordered state of correlated metal $\text{SrVO}_{3}$. Numerical simulations utilizing a classical tensor network implementation of variational quantum eigensolvers allow us to simulate large models with up to $54$ qubits and encompassing up to four bands in the correlated subspace, which is indicative of the complexity that our framework can address. Through these methods we demonstrate the potential of classical pre-optimization and downfolding techniques for enabling efficient materials simulation using quantum algorithms.
@@ I Introduction
The full many-body Hamiltonian describing a general material includes $M^{4}$ terms, where $M$ the number of orbitals. For realistic systems this can be prohibitively large. A promising computational method for generating a compressed representation of materials from first principles is ab initio downfolding [11, 12, 13, 14]. This technique utilizes a DFT starting point in order to derive a material-specific many-body Hamiltonian within a strongly-correlated active space of interest, which can be subsequently solved using a range of techniques, including exact diagonalization [15, 16, 17, 18], quantum Monte Carlo [19, 20, 21], diagrammatic approximations [22], and beyond. The many-body Hamiltonian resulting from downfolding and describing the physics of the strongly-correlated region is typically of the extended Hubbard form [23]:
@@
where $\sigma,\rho$ are spin indices, $\mathbf{R},\mathbf{R}^{\prime}$ denote lattice vectors, and $i,j$ run over the electronic bands of the system. Downfolding methods have been successfully applied to the description of charge-ordered systems and charge density waves [24, 25], high-temperature superconductors [13, 26, 27, 28, 29], and beyond. However, the unfavorable scaling of classical methods for obtaining eigenstates of the Hamiltonian of Eq. (I) has thus far prevented downfolding methods from being used for large systems, which can result in finite-size effects [30, 31, 32], and limits the number of bands which may realistically be included in these models.
@@
Quantum computers are a promising technology for the simulation of many-body quantum systems [33, 34], especially with regards to the number of qubits needed to simulate Hamiltonians of the form of eq. (I), as many algorithms can utilize a 1:1 correspondence between the qubits and the spin-orbitals of the system. This constitutes the simulation of strongly-correlated materials and Hubbard-like models an ideal problem for solution on quantum hardware [35, 36], and the performance of popular variational quantum algorithms in terms of producing an accurate ground state has recently been the subject of a detailed benchmark [37]. The potential of emerging fault-tolerant quantum computers to obtain properties of extended Hubbard models as that of Eq. (I) has been discussed in Ref. [38], where the necessary resources to access experimentally-relevant quantities have been estimated.
@@
The utility of quantum computers for materials simulation is currently limited by the fact that, as previously mentioned, the number of terms of the full Hamiltonian of a typical material scales as $M^{4}$. This can in turn lead to extremely high resource estimates in terms of gates and run times. Here we circumvent this limitation by demonstrating that current and near-term quantum algorithms can produce quantitatively and qualitatively accurate results for ground states of strongly-correlated materials, at a modest computational cost. We achieve this by utilizing compressed representations obtained via ab initio downfolding, and classical tensor networks simulations of the variational quantum eigensolver (VQE) [39]. We show that our VQE energy for the ground state is quantitatively accurate when compared to results obtained within density matrix renormalization group (DMRG), and that we obtain the correct behavior for the ground state wavefunction and correlation functions across different scenarios, and specifically the antiferromagnetic behavior of the quasi-1D cuprate $\text{Ca}_{2}\text{CuO}_{3}$ [40], the excitonic insulating ground state of the two-dimensional material $\text{WTe}_{2}$ [41, 7], and the charge-ordered state of the correlated metal $\text{SrVO}_{3}$ [42, 43]. These results, combined with demonstrations that quantum simulation of downfolded Hamiltonians can yield accurate excited state properties in strongly correlated molecular systems [44, 45, 46], highlight the strong potential of current and emerging quantum computers for simulating the properties of strongly-correlated materials at a modest computational cost, when combined with ab initio downfolding.
Figure 1: Schematic illustration of ab initio downfolding. Starting from a low-level description of the full system of interest, an active space is identified, and a many-body lattice Hamiltonian is generated, representing the physics within the active space. A typical form for the downfolded Hamiltonian is the extended Hubbard form of eq. (I), where each lattice site can be associated with different hopping and Coulomb terms.
@@ II.1 Ab initio downfolding
The aim of ab initio downfolding approaches is to generate a many-body Hamiltonian on a lattice, representing an active space of interest for a given material, typically within the low-energy region in the vicinity of the Fermi level [11, 12, 13, 23]. This is schematically illustrated in Fig. 1. Downfolding often utilizes a first-principles description such as DFT as a starting point. An exchange term may be included in downfolded Hamiltonians [23], however we will ignore it here given its small magnitude for our studied systems, consistent with previous works [24]. In order to describe periodic solids as a lattice model, we work within the basis of maximally-localized Wannier functions [47], which we will denote as $\phi$. Here the Wannier functions are centered at the different lattice sites $\mathbf{R}$ appearing in Eq. (I), with $i,j$ indexing the $i^{\text{th}}$ and $j^{\text{th}}$ Wannier function respectively. Starting from a Kohn-Sham DFT calculation performed within Quantum Espresso [48], the Wannier representation of the studied systems within the active space of interest is obtained using the Wannier90 code [49]. The Wannierization procedure yields the hopping terms appearing in Eq. (I) as
@@
where $H_{KS}$ the Kohn-Sham Hamiltonian and $V$ the volume of the unit cell. For $\mathbf{R}=\mathbf{R}^{\prime}$ this term represents the on-site potential of site $\mathbf{R}$. Moreover, following the wannierization using Wannier90, we utilize the wan2respack [50] and RESPACK [23] software packages in order to obtain the two-body Coulomb integrals appearing in Eq. (I) as
@@ II.2 Tensor network VQE simulation
Unless otherwise explicitly stated, we use the maximum bond dimension $\chi_{\rm max}=2^{n_{q}/2}$, where $n_{q}$ the number of qubits. Since this bond dimension is sufficient to exactly represent an arbitrary wavefunction on $n_{q}$ qubits, the MPS representation yields the energy expectation value with respect to the exact PQC. For all tensor operations in this work we have used the ITensor software package [58]. We provide details on our optimization procedure in Section A of the Appendix. In preparing our PQC, we utilize two ansätze in this work, a number-preserving (NP) ansatz [36], which was designed specifically within the context of solving single-band Hubbard models, and a more generic excitation-preserving (EP) [59] ansatz, which allows us to straightforwardly prepare our circuits in the case of multi-band Hubbard models. The performance of both ansätze for producing accurate ground states has been discussed in Ref. [37]. We apply the ansätze to the solution of the non-interacting ($U=0$) Hubbard representation of our systems, prior to which we apply $R_{z}(\theta)$ gates to each qubit, which improves the optimization procedure.
II.3 Hamiltonian compression and the measurement problem
Our aim is to obtain the ground state of the Hamiltonian of Eq. (I), for studied materials with an active space represented by different sets of $U_{i\mathbf{R}j\mathbf{R}^{\prime}}$ and $t_{i\mathbf{R}j\mathbf{R}^{\prime}}$ matrices. Before presenting our results, it is worth discussing some aspects of the complexity of this problem. For a Hubbard model on an $N_{x}\times N_{y}$ square lattice, with $N_{b}$ electronic bands included in the active space of the studied system, $n_{q}=2N_{x}N_{y}N_{b}$ is the number of qubits needed to simulate its properties, where the factor $2$ accounts for spin. Denoting $N=N_{x}N_{y}$, the number of terms appearing in the Hamiltonian of Eq. (I) is classified as follows, considering only nearest-neighbor contributions:
  • On-site potential $t_{i\mathbf{R}i\mathbf{R}}$: For the on-site energies, there are $2\times N_{b}\times N$ terms.
  • Total number of terms:
    $n_{\text{terms}}=N_{b}\times\bigg{[}(N_{b}+2)[N_{x}\times(N_{y}-1)+$
    $N_{y}\times(N_{x}-1)]+3N_{b}N\bigg{]}$
We do not include inter-band hopping between different sites to ensure a manageable computational cost for the systems studied here. Boundary effects have been implicitly considered by restricting hopping and off-site interactions to existing nearest-neighbor pairs within the lattice. It becomes clear that the compressed representation arising from ab initio downfolding, leads to a much improved scaling of $N_{b}^{2}N_{x}N_{y}$ to leading order, compared to the $(N_{b,f}N_{x}N_{y})^{4}$ scaling of the full many-body Hamiltonian. Note that here $N_{b,f}$ is the full number of bands of the system, and not only the ones within an active space, and the total number of terms appearing in the full-many body Hamiltonian is equal to $2(N_{x}N_{y}N_{b,f})^{2}+4N_{x}N_{y}N_{b,f})^{4}$. Therefore, the compression scales as $N_{b}^{2}/(N_{b,f}^{4}N_{x}^{3}N_{y}^{3})$. For the specific examples we study below, we precisely quantify the degree of Hamiltonian compression achieved through our downfolding procedure. The number of bands $N_{b,f}$ entering the full many-body Hamiltonian is set to be equal to the number of occupied bands, plus any empty states that enter the active space of the material within the downfolded representation. To also enable a more fair comparison, we give the number of terms of the full many-body Hamiltonian when it is restricted within the active space of the compressed representation.
The compressed representation of materials becomes particularly important when considering the problem of measuring expectation values of their Hamiltonian on quantum hardware. For a system described my $M$ qubits, depending on the specific measurement strategy, such as measuring qubit-wise commuting terms simultaneously, measuring a family of non-crossing (NC) terms simultaneously, or performing a basis rotation (BR) grouping, the lower bound for the number of measurements required in order to determine the energy for a single VQE energy evaluation scales as $M^{2}-M^{6}$ [60, 61]. Although a more detailed investigation is warranted, these lower bounds suggest that it is conceivable, when using methods such as NC and BR groupings with an idealized scaling close to $M^{2}$ for the number of measurements, to achieve accurate simulations of strongly-correlated materials on near-term quantum hardware. Even for less favorable scaling in real devices, the compressed representations of our systems (no more than $M=54$ qubits) could be particularly important in this direction.
Finally, it is worth highlighting that when performing a finite amount of measurements for different observables on real quantum hardware, the expectation values necessarily have a statistical uncertainty associated with them. Our classical tensor network simulation of VQE does not currently include shot noise as part of the simulation, but rather it represents the ultimate accuracy that could be expected from the simulation of the compressed material Hamiltonians.
II.4 Downfolded Hamiltonian simulation with near-term and fault-tolerant hardware
It is worth emphasizing that the quantum simulation of materials using representations as the Hamiltonian of Eq. (I) has been the subject of detailed benchmarks, and found to likely be feasible with near-term resources, without necessitating fully fault-tolerant quantum computers [61]. Indeed, as we outline in Section III for the different materials studied here, the number of qubits and two-qubit gates are well within the thresholds set in Ref. [61] as what might be feasible on near-term hardware. Moreover, as also detailed in Section III, by assuming that the fidelity of two-qubit gates is $99.9$[62], we find that for all studied systems here our circuit fidelities are in the range of $52-75$% thanks to our compression of the Hamiltonian, which places them well above the $10$% threshold that has been identified for obtaining useful results [63]. We emphasize that while this is only meant to provide a rough estimate, it clearly demonstrates the importance of the compression for reducing the measurement problem.
@@ II.4 Downfolded Hamiltonian simulation with near-term and fault-tolerant hardware
Nevertheless, as we aim to build increasing complexity into the reduced Hamiltonians obtained through downfolding or related methods, and approach the full many-body limit, it will become necessary to utilize fault-tolerant architectures [38]. However, our work here is mostly agnostic to this. We can obtain a rough estimate of the resources that would be needed for a fault-tolerant simulation of the downfolded Hamiltonians using the ansatz considered here, through computing different metrics. For example, one may compute the 1-norm of the Hamiltonian of eq. (I):
$||H||_{1}=\sum_{ij}\sum_{\mathbf{R}\mathbf{R}^{\prime}}|\tilde{t}_{i\mathbf{R}j\mathbf{R}^{\prime}}|+\frac{1}{2}\sum_{ij}\sum_{\mathbf{R}\mathbf{R}^{\prime}}|\tilde{U}_{i\mathbf{R}j\mathbf{R}^{\prime}}|,$ (7)
where $\tilde{t},\tilde{U}$ are dimensionless, obtained through dividing the hopping and Coulomb matrices by the value of the dominant intra-band nearest-neighbor hopping. Our definition for the 1-norm includes summations over all neighbors, despite the fact that within our VQE simulations we only include nearest neighbor terms. The localized nature of the Wannier functions results in terms beyond nearest neighbor being small, and our values for the 1-norm should be considered as an upper bound for the Hamiltonians we actually simulate. The 1-norm provides a bound for the Trotter error [64], and smaller values for this quantity are associated with more efficient quantum phase estimation [65].
@@
For the purpose of fault-tolerant simulation it is also interesting to estimate the number of $T$-gates which will be necessary in order to implement our circuits, where the one-qubit and two-qubit operators are approximated with a required accuracy $\epsilon$. As described previously, we apply an $R_{z}$ gate to each qubit prior to the application of the unitary operators associated with the variational ansatz. The approximation of a one-qubit gate has been found to be possible with $1.15\log_{2}(1/\epsilon)$ $T$-gates [66]. From this we have $n_{q}\times 1.15\log_{2}(1/\epsilon)$ $T$-gates associated with this step. Additionally, both the NP and EP ansätze involve the application of two-qubit gates with two parameters each. For a total of $n_{\text{params}}$ parameters in the variational ansatz we have $n_{\text{params}}/2$ two-qubit gates, each of which requires $\mathcal{O}[16\log(1/\epsilon)+32]$ $T$-gates to represent [67]. Therefore, we have a total of $n_{q}\times 1.15\log_{2}(1/\epsilon)n+\frac{n_{\text{params}}}{2}\times\mathcal{O}[16\log(1/\epsilon)+32]$ $T$-gates required for the fault-tolerant simulation of our systems with accuracy $\epsilon$.
@@ III Results
$\text{Ca}_{2}\text{CuO}_{3}$ $10\times 1$ $1$ $6.005$ $6.028$ $99.3\%$
$\text{WTe}_{2}$ $2\times 2$ $4$ $115.029$ $115.097$ $96.2\%$
$\text{SrVO}_{3}$ $3\times 3$ $3$ $-105.383$ $-105.365$ $31.8\%$
Table 1: Studied systems, simulated lattice sizes, number of bands in the subspace, DMRG energies, and best VQE energies and fidelities with respect to the DMRG solution. As outlined in the main text, we solve the Hubbard model representation of $\text{Ca}_{2}\text{CuO}_{3}$ at half-filling, of $\text{WTe}_{2}$ at full filling of the valence bands, and of $\text{SrVO}_{3}$ at half-filling of the lowest band.
system $n_{q}$ $n_{2q,G}$ circuit fidelity $||H||_{1}$ $n_{\text{terms}}$
$\text{Ca}_{2}\text{CuO}_{3}$ $20$ $290$ $74.8$% $2.67\times 10^{2}$ $37$
$\text{WTe}_{2}$ $32$ $652$ $52.1$% $3.31\times 10^{2}$ $288$
$\text{SrVO}_{3}$ $54$ $484$ $55.8$% $2.315\times 10^{3}$ $423$
Table 2: Parameters relating to the near-term and fault-tolerant quantum simulation of the studied systems. Specifcally, we give the number of qubits ($n_{q}$) required to simulate the downfolded Hamiltonians, the number of two-qubit gates $n_{2q,G}$ in our circuits, and the respective circuit fidelity if one assumes a $99.9\%$ fidelity for the individual gates, as well as the 1-norm and the number of terms of the downfolded Hamiltonians.
Fig. 2 visualizes the structures of the systems we study here. These are chosen to demonstrate the capability of our approach to correctly predict the ground state properties of diverse strongly-correlated materials; $\text{Ca}_{2}\text{CuO}_{3}$ is a quasi-1D cuprate known to display antiferromagnetic behavior along the Cu-O chains [40]; $\text{WTe}_{2}$ has been proposed to host an excitonic ground state, i.e., one where correlated electron-hole pairs form spontaneously [41, 7]; and $\text{SrVO}_{3}$ is a correlated metal that exhibits substantial charge ordering [42, 43]. Table 1 summarizes the number of bands included in the active space of each system, the lattice size on which the downfolded Hamiltonian is solved, the DMRG energy, and the best value for the energy and fidelity obtained from the ten independent VQE optimizations we perform. For $\text{Ca}_{2}\text{CuO}_{3}$ we only perform a single optimization, as the simpler energy landscape of this system makes this sufficient for finding a low-energy, high-fidelity solution. More details on the optimization for each system are given below. In Section A of the Appendix we also give the VQE energy and fidelity values for all optimizations performed for $\text{WTe}_{2}$ and $\text{SrVO}_{3}$. Moreover, we summarize details of all DFT and cRPA calculations employed for these systems in Section B of the Appendix. We now discuss the properties of the ground state wavefunctions of the materials of Fig. 2 as obtained within our combined ab initio downfolding/VQE approach.